### v1

@@ 1. Introduction
In optimization theory, when studying the limit of a sequence of optimization problems, one seeks a natural notion of convergence of the data that guarantees convergence of both optimal values and minimizers. In other words, a form of variational convergence that allows one to pass to the limit in variational problems. The first concept of this kind was introduced by Wijsman in finite dimensions under the name infimal convergence [78, 79], motivated by questions in statistical decision theory [31]. In the 1970s, De Giorgi and Franzoni extended this perspective to general topological spaces under the name $\Gamma$-convergence [47], which has since become one of the central tools of the calculus of variations.
@@
A further milestone was reached by Attouch in 1977, who proved that Mosco convergence of convex functions is equivalent to the convergence of their subdifferentials in the sense of Painlevé–Kuratowski set convergence, up to a normalization constant [3]. This fundamental result, now known as Attouch’s theorem, has become a cornerstone of variational analysis. It has found numerous applications: to stability in optimization and variational problems [14, 13, 9, 10, 11, 12], to the analysis of parabolic variational inequalities [6], to numerical stability in optimization algorithms [51], and to structural results such as generalized second-order derivatives of convex functions [68, 67, 34, 33, 52, 63] and differentiability of Lipschitz set-valued mappings [39]. Attouch’s theorem also remains influential in applied areas such as nonsmooth mechanics and the study of subgradient evolution systems [5].
@@ 3.1. Notations and basic definitions
Finally, $X$ is said to be a weakly compactly generated (WCG) space if there exists a weakly compact set $A\subset X$ such that the linear span of $A$ is dense in $X$ with respect to the norm topology. In this definition, $A$ may be assumed, without loss of generality, to be convex and symmetric (see, e.g., [43, Chapter 13]).
@@
$\partial f:=\{(x,x^{*})\in X\times X^{*}:x^{*}\in\partial f(x)\}.$
@@ 4. Limits of convex functions and subdifferentials
Let $X$ be a WCG Banach space, and let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be a sequence of lower semcontinuous proper convex functions such that the sequence $(f_{n}^{*})_{n\in\mathbb{N}}$ satisfies the equicoercivity property
@@
Consider the sequence of functions $f_{n}(x):=\tfrac{1}{n}|x|+n$. On the one hand, we have
@@
Now, let $x^{\ast}\in C\setminus\operatorname{Ls}_{w^{\ast}}C_{n}$. On the one hand, we have $f^{\ast}(x^{\ast})=0$. On the other hand, suppose that there exists a sequence $x_{n}^{\ast}\to x^{\ast}$ with $\limsup_{n}f^{\ast}_{n}(x^{\ast}_{n})<\infty$. Then the inclusion $x_{n}^{\ast}\in C_{n}$ must hold for all sufficiently large $n\in\mathbb{N}$; otherwise, we would obtain $\limsup_{n}f^{\ast}_{n}(x^{\ast}_{n})=\infty$. However, this would imply $x^{\ast}\in\operatorname{Ls}_{w^{\ast}}C_{n}$, which is a contradiction. Therefore, the only possible alternative is
@@
It is worth noting that the equicoercivity property given in (4.1) can be readily verified under the stronger assumption of uniform coercivity, that is, there exist constants $\alpha>0$ and $\beta\in\mathbb{R}$ such that
@@
By taking the Fenchel conjugate in (4.2) (see, e.g., [28, Fact 4.4.9]), one obtains that this condition is equivalent to the uniform boundedness of the functions in a neighborhood of the origin, which menas, there exist $\rho,\varepsilon>0$ such that
@@ 5. Some Preliminary Results
Since the sequences $(f_{n}(a_{n}))_{n\in\mathbb{N}}$ and $(\left\lVert y_{n}^{\ast}\right\rVert\left\lVert a_{n}\right\rVert)_{n\in\mathbb{N}}$ are bounded above, then we get
@@
Let us prove that $x_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}$. Since $(x_{n}^{\ast})_{n\in\mathbb{N}}$ is norm bounded and $\{x_{j}:j\in\mathbb{N}\}$ is dense in $X$, all we need to proof is that
@@ 6.2. Separable reduction and rich families
We say that a property $\mathcal{P}$ conecerning a Banach space $X$, is separably reducible (or separably determined) if there exists a family $\mathcal{F}$ of separable subspaces of $X$ such that $\cup\left\{F:F\in\mathcal{F}\right\}=X$ and the property $\mathcal{P}$ holds in $X$ if and only if for every $F\in\mathcal{F}$ the property $\mathcal{P}$ holds at $F$.
Separable reducibility is a common tool in functional analysis, as it often leads to substantial simplifications and, in some cases, provides the only viable approach to proving certain results. A simple example of a separably reducible property is continuity of a mapping. For a list of some less trivial properties that are separable determined in analysis, we refer the reader to [44]. There are special families of separable subspaces of a given Banach space that allow to implement a separable reduction, that is, the possibility to “lift” a property from separable subspaces to the entire space, since they allow dealing with countable many different properties. These special families are called rich families, and they were introduced by J.M Borwein and W.B Moors in [26]. Rich families were deeply discussed in connection with a number of properties related to differentiability and subdifferentiability in [53] and [45].
@@
The main property and the power is showed in the following fundamental fact (see [26, Proposition 1.1]).
@@ 6.4. Preliminary Results on Separable Reduction Techniques
The first result concerns the nonemptiness of the domain of the upper epilimt, in order to show that the hyphoteses ${\rm dom}\kern 1.19995pt(\Gamma\text{-}\displaystyle\limsup_{n}f_{n}))\neq\emptyset$ of Theorem 4.1 is true if we restrinct each $f_{n}$ to certain separable subspaces.
@@
Let $X$ be a WCG Banach space. Let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be a sequence of proper lower semicontinuous convex function satisfying the equicoercivity property (4.1). Then there exists a rectangle rich family $\mathcal{R}\subset\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ such that for every $V\times Y\in\mathcal{R}$ and every sequence $(v^{\ast}_{n})_{n\in\mathbb{N}}\subset V^{\ast}$
@@
On the one hand, by the property of $\mathcal{F}$, there exists $y^{\ast}_{n}\in Y$ such that $y^{\ast}_{n}|_{V}=v^{\ast}_{n}$ and $\|v^{\ast}_{n}\|=\|y^{\ast}_{n}\|$. On the other hand, by the properties of each $\mathcal{R}_{n}$, we have
@@
Denote $f:=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$ and fix $x^{\ast}\in X^{\ast}$. Let $\gamma:=V\times Y\in\mathcal{R}$ be such that $x^{\ast}\in Y$. By Lemma 6.1, we may restrict to the case $f^{\ast}(x^{\ast})<\infty$.
@@
and the the normalization condition (N.C.) holds, that is, there exists $(a,a^{*})\in\partial f$ and a sequence $(a_{n},a_{n}^{*})\in\partial f_{n}$ such that $a_{n}\to a,\;a_{n}^{*}\overset{\ast}{\rightharpoonup}a^{*}\text{ and }f_{n}(a_{n})\to f(a).$

### v2

@@ 1. Introduction
In optimization theory, when studying the limit of a sequence of optimization problems, one seeks a natural notion of convergence of the data that guarantees convergence of both optimal values and minimizers; in other words, a form of variational convergence that allows one to pass to the limit in variational problems. The first concept of this kind was introduced by Wijsman in finite dimensions under the name infimal convergence [78, 79], motivated by questions in statistical decision theory [31]. In the 1970s, De Giorgi and Franzoni extended this perspective to general topological spaces under the name $\Gamma$-convergence [47], which has since become one of the central tools of the calculus of variations.
@@
A further milestone was reached by Attouch in 1977, who proved that Mosco convergence of convex functions is equivalent to the convergence of their subdifferentials in the sense of Painlevé–Kuratowski set convergence, up to a normalization constant [3]. This fundamental result, now known as Attouch theorem, has become a cornerstone of variational analysis. It’s applications are numerous: to stability in optimization and variational problems [14, 13, 9, 10, 11, 12], to the analysis of parabolic variational inequalities [6], to numerical stability in optimization algorithms [51], and to structural results such as generalized second-order derivatives of convex functions [68, 67, 34, 33, 52, 63] and differentiability of Lipschitz set-valued mappings [39]. Attouch’s theorem also remains influential in applied areas such as nonsmooth mechanics and the study of subgradient evolution systems [5].
@@ 3.1. Notations and basic definitions
Finally, $X$ is said to be a weakly compactly generated (WCG) space if there exists a weakly compact set $A\subset X$ such that the linear span of $A$ is dense in $X$ with respect to the norm topology. In this definition, $A$ may be assumed, without loss of generality, to be convex and symmetric (see, e.g., [43, Chapter 13]). For some examples of $WCG$ spaces, we refer the reader to [43, Chapter 13 and Theorem 14.9].
@@
$\partial f:=\{(x,x^{*})\in X\times X^{*}:x^{*}\in\partial f(x)\},$
@@ 4. Limits of convex functions and subdifferentials
Let $X$ be a WCG Banach space, and let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be a sequence of lower semicontinuous proper convex functions such that the sequence $(f_{n}^{*})_{n\in\mathbb{N}}$ satisfies the equicoercivity property
@@
Consider the sequence of functions $f_{n}(x):=\tfrac{1}{n}|x|+n$. On one hand, we have
@@
Now, let $x^{\ast}\in C\setminus\operatorname{Ls}_{w^{\ast}}C_{n}$. On one hand, we have $f^{\ast}(x^{\ast})=0$. On the other hand, suppose that there exists a sequence $x_{n}^{\ast}\to x^{\ast}$ with $\limsup_{n}f^{\ast}_{n}(x^{\ast}_{n})<\infty$. Then the inclusion $x_{n}^{\ast}\in C_{n}$ must hold for all sufficiently large $n\in\mathbb{N}$; otherwise, we would obtain $\limsup_{n}f^{\ast}_{n}(x^{\ast}_{n})=\infty$. However, this would imply $x^{\ast}\in\operatorname{Ls}_{w^{\ast}}C_{n}$, which is a contradiction. Therefore, the only possible alternative is
@@
It is worth noting that the equicoercivity property given in (4.1) is true when assuming that for each $t\in\mathbb{R}$ there is a $w^{\ast}$-compact set $K_{t}\subset X^{\ast}$ such that for every $n\in\mathbb{N}$, $\{x^{\ast}\in X^{\ast}:f^{\ast}_{n}(x^{\ast})\leq t\}\subset K_{t}$, which is a standard hypothesis to assure convergence of minima in $\Gamma$-convergence. Also, it can be readily verified under the stronger assumption of uniform coercivity, that is, there exist constants $\alpha>0$ and $\beta\in\mathbb{R}$ such that
@@
By taking the Fenchel conjugate in (4.2) (see, e.g., [28, Fact 4.4.9]), one obtains that this condition is equivalent to the uniform boundedness of the functions in a neighborhood of the origin, which means, there exist $\rho,\varepsilon>0$ such that
@@ 5. Some Preliminary Results
Since the sequences $(f_{n}(a_{n}))_{n\in\mathbb{N}}$ and $(\left\lVert y_{n}^{\ast}\right\rVert\left\lVert a_{n}\right\rVert)_{n\in\mathbb{N}}$ are bounded above, we then get
@@
Let us prove that $x_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}$. Since $(x_{n}^{\ast})_{n\in\mathbb{N}}$ is norm bounded and $\{x_{j}:j\in\mathbb{N}\}$ is dense in $X$, all we need to prove is that
@@ 6.2. Separable reduction and rich families
We say that a property $\mathcal{P}$ concerning a Banach space $X$, is separably reducible (or separably determined) if there exists a family $\mathcal{F}$ of separable subspaces of $X$ such that $\cup\left\{F:F\in\mathcal{F}\right\}=X$ and the property $\mathcal{P}$ holds in $X$ if and only if for every $F\in\mathcal{F}$ the property $\mathcal{P}$ holds at $F$.
Separable reducibility is a common tool in functional analysis, as it often leads to substantial simplifications and, in some cases, provides the only viable approach to proving certain results. A simple example of a separably reducible property is continuity of a mapping. For a list of some less trivial properties that are separable determined in analysis, we refer the reader to [44]. There are special families of separable subspaces of a given Banach space that allow the implementation of a separable reduction, that is, the possibility to “lift” a property from separable subspaces to the entire space, since they permit countable many different properties to be dealt with. These special families are called rich families, and they were introduced by J.M Borwein and W.B Moors in [26]. Rich families were deeply discussed in connection with a number of properties related to differentiability and subdifferentiability in [53] and [45].
@@
The main property and the power of rich families is shown in the following fundamental fact (see [26, Proposition 1.1]).
@@ 6.4. Preliminary Results on Separable Reduction Techniques
The first result concerns the nonemptiness of the domain of the upper epilimit, in order to show that the hyphotesis ${\rm dom}\kern 1.19995pt(\Gamma\text{-}\displaystyle\limsup_{n}f_{n}))\neq\emptyset$ of Theorem 4.1 is true if we restrict each $f_{n}$ to certain separable subspaces.
@@
Let $X$ be a WCG Banach space. Let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be a sequence of proper lower semicontinuous convex functions satisfying the equicoercivity property (4.1). Then there exists a rectangle rich family $\mathcal{R}\subset\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ such that for every $V\times Y\in\mathcal{R}$ and every sequence $(v^{\ast}_{n})_{n\in\mathbb{N}}\subset V^{\ast}$
@@
On one hand, by the property of $\mathcal{F}$, there exists $y^{\ast}_{n}\in Y$ such that $y^{\ast}_{n}|_{V}=v^{\ast}_{n}$ and $\|v^{\ast}_{n}\|=\|y^{\ast}_{n}\|$. On the other hand, by the properties of each $\mathcal{R}_{n}$, we have
@@
Denote $f:=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$ and fix $x^{\ast}\in X^{\ast}$. Let $\gamma:=V\times Y\in\mathcal{R}$ be such that $x^{\ast}\in Y$. By Lemma 6.1, we may assume that $f^{\ast}(x^{\ast})<\infty$.
@@
and the normalization condition (N.C.) holds, in other words, there exists $(a,a^{*})\in\partial f$ and a sequence $(a_{n},a_{n}^{*})\in\partial f_{n}$ such that $a_{n}\to a,\;a_{n}^{*}\overset{\ast}{\rightharpoonup}a^{*}\text{ and }f_{n}(a_{n})\to f(a).$